OpenStudy (anonymous):

how do you find the derivative of (4)/(7x) so that it may equals (-4)/(7x^2). please show calc involved

OpenStudy (phi):

rewrite it using a negative 1 exponent \[ \frac{4}{7} x^{-1} \] now use the power rule

OpenStudy (phi):

do you know the power rule? \[ \frac{d}{dx} x^n = n x^{n-1} \]

OpenStudy (anonymous):

no i dont